# Can Emulating Semantic Translation Help LLMs with Code Translation? A Study Based on Pseudocode

### v2

@@ Abstract.
Large language models (LLMs) show great potential in code translation. However, accurate translation remains challenging when using the commonly adopted direct code-to-code translation approach, which converts a program into the target programming language (PL) in a single step. Inspired by the success of incorporating intermediate steps to guide LLMs in resolving challenging tasks, we explore pseudocode-based code translation, which emulates the human semantic translation by first interpreting the program’s intent and logic into pseudocode and then implementing it in the target PL. We find that pseudocode-based translation helps translate programs that direct translation struggles to handle. Nonetheless, the effectiveness, advantages, and limitations of this approach remain underexplored. To bridge this gap, we present an empirical study on pseudocode-based code translation, aiming to investigate its effectiveness in enhancing the direct translation approach, illuminate its effective usage, and identify limitations hindering its potential benefits. By comparing direct and pseudocode-based translation approaches on 9,690 translation tasks across six PLs with five popular LLMs, we demonstrate that pseudocode-based translation can effectively complement direct translation, particularly when translating from flexible to rigid PLs or dealing with low-resource Rust. Based on these findings, we suggest adopting strategies that combine the complementary strengths of both approaches to enhance code translation accuracy. We also reveal the advantages of pseudocode-based translation in disentangling translations of complicated programs and mitigating distractions from detailed implementations in original programs, as well as its limitations due to incorrect, incomplete, or ambiguous pseudocode.
@@ 1. Introduction
Code translation, also known as transpilation, refers to the automatic conversion of a program written in one programming language (PL) to another while preserving its semantics (i.e., program functionality) (Nguyen et al., 2013; Rozière et al., 2020; Pan et al., 2024). With the rapid evolution of PLs and the diverse requirements of software applications, code translation has gained significant attention in both research and industry due to its wide range of practical applications. For example, code translation facilitates the migration of legacy software systems built based on obsolete PLs to modern PLs (Echeverria et al., 2015; Kalia et al., 2021a, b). It also facilitates efficient prototyping and development across multiple PLs for multi-platform software (Liu et al., 2023; Nguyen et al., 2015; Wu et al., 2010). However, achieving accurate automated code translation remains a challenging task due to the inherent differences in syntax and features among various PLs (Pan et al., 2024).
@@
Over the past few decades, various approaches have been proposed to automate code translation and address the associated challenges (Chen et al., 2025; Jiao et al., 2023). Earlier methods primarily relied on statistical or neural machine translation (Nguyen et al., 2013; Chen et al., 2018; Rozière et al., 2020, 2022; Liu et al., 2023), while recent practices start leveraging the powerful large language models (LLMs) as code translators (Pan et al., 2024; Yang et al., 2024). Nevertheless, existing works mainly use a direct code-to-code translation approach, where the original program is taken as input to generate the translated program in a single step. In this step, LLMs may perform multiple tasks implicitly, such as understanding the semantics of the original program and generating the target program in another PL. Studies have shown that such direct code-to-code translation remains challenging, especially across PLs with significant differences in syntax and features (Pan et al., 2024; Macedo et al., 2025).
@@
Decomposing challenging tasks into intermediate steps has proven effective in guiding LLMs to emulate successful human workflows in various coding tasks (e.g., splitting planning and implementation for code generation (Dong et al., 2024; Li et al., 2025) and fault localization and patch generation for debugging (Xia et al., 2024)). Notably, we observe that code translation can also benefit from explicitly implementing intermediate steps that emulate human practices in translation. Specifically, human translators often perform semantic translation (Newmark, 1981) based on the meaning of the given text, first distilling the intent into a language-agnostic meaning and then rendering it in the target language (Newmark, 1981). Meanwhile, a recent study demonstrates that pseudocode, as a PL-agnostic representation of code intent and logic widely used in textbooks and research papers, can effectively guide code generation across PLs (Wu et al., 2025). Based on these insights, we further recognize that emulating semantic translation for code based on pseudocode can be a solution for the tasks that direct translation struggles with. For example, when translating a simple C++ program (Figure 1(a)) to Rust, all five studied LLMs (e.g., Qwen2.5-Coder-32B-Instruct) failed to produce an accurate translation via a one-step direct translation (Figure 1(b)). In comparison, following a clear pseudocode (Figure 1(c)) summarized from the original C++ program, these LLMs can successfully implement the original program’s functionality in Rust (Figure 1(d)). However, introducing explicit intermediate steps also has limitations. For example, it may introduce noise or suffer from semantic loss during information transmission (Liu et al., [n. d.]), which may conversely hinder accurate code translation. In general, the effectiveness, advantages, and limitations of such pseudocode-based code translation compared to the commonly adopted direct translation remain unclear.
@@
In this work, we conduct the first empirical study on pseudocode-based code translation to bridge the gap in understanding this under-explored approach. Specifically, we assessed the performance of five translation strategies: the widely-adopted direct translation approach, two pseudocode-based strategies, and two hybrid strategies that combine direct and pseudocode-based translation. Through four research questions (RQs), we investigate the overall effectiveness of pseudocode-based code translation compared with the widely-adopted direct translation (RQ1), the helpfulness of pseudocode for varying PL pairs (RQ2), the advantages of pseudocode as an intermediary compared to a concrete PL, which also proved helpful in code translation (Macedo et al., 2025) (RQ3), and how the quality of pseudocode affects translation outcomes (RQ4). We also examine the successes and failures of pseudocode-based code translation via case studies and discuss future directions for further harnessing its potential benefits. The experiment is conducted with six popular PLs on solution programs of 323 LeetCode problems in three difficulty levels in LiveCodeBench (Jain et al., 2025), resulting in 9,690 translation tasks for each of five studied LLMs.
@@
Our experimental results reveal several interesting findings and actionable insights. Specifically, we identify the complementary role of pseudocode-based code translation in effectively handling code translation tasks that direct translation struggles with. By combining the results of direct translation and pseudocode-based translation, the studied LLMs achieve average improvements in pass@10 rates over direct translation alone of 4.10%, 7.44%, and 13.75% on easy-, medium-, and hard-level problems, respectively. We also find that pseudocode-based translation can benefit varying source-target PL pairs, with more substantial helpfulness for translation from flexible PLs (e.g., Python) to more rigid PLs (e.g., Go) or involving the low-resource PL Rust. These findings suggest the adoption of a hybrid strategy combining direct and pseudocode-based translation results with further test-based selection to leverage the complementary advantages of both approaches, in particular when translating from flexible to rigid PLs or dealing with low-resource PLs. Moreover, with higher-quality pseudocode, we observed that the studied LLMs show the potential to achieve pass rates of 0.9646–0.9835, 0.8861–0.9512, and 0.6747–0.8286 on three-level tasks, respectively. The identified bottlenecks of pseudocode-based code translation stem from both code understanding and generation capabilities of LLMs. Our case studies also highlight the advantages of pseudocode-based translation in disentangling understanding and generation on complicated programs and mitigating distractions from the PL-specific details in original programs, as well as its limitations caused by incorrect, incomplete, or ambiguous pseudocode. These findings further illuminate the effective use of pseudocode-based translation and inspire future research on enhancing code translation through pseudocode.
@@
Figure 1. An Example of Correct C++-to-Rust Translation by Qwen2.5-Coder-32B-Instruct based on Pseudocode (The other four studied LLMs show similar symptoms in this case.)
@@
  • We conduct the first empirical study on pseudocode-based code translation, exploring the effectiveness, advantages, and limitations of explicit semantic code translation via pseudocode in improving code translation accuracy by emulating human translation practices.
  • We compare four pseudocode-based translation strategies with direct translation across 9,690 translation tasks based on programs in six popular PLs. With five popular LLMs as the code translator, we systematically investigate the effectiveness of pseudocode-based code translation in enhancing direct translation, its helpfulness for varying PL pairs, and its advantages and limitations, aiming to provide insights for its effective use and inspire future enhancement.
  • Our study reveals interesting findings, including the complementary role of pseudocode-based translation in enhancing direct code translation, particularly for flexible-to-rigid PL translation and for low-resource Rust, and the potential and bottlenecks of pseudocode-based code translation. These findings yield actionable insights, including the adoption of a hybrid strategy combining direct and pseudocode-based translation results to enhance code translation accuracy, as well as the space in pseudocode quality to further harness the benefits of the approach.
The rest of the paper is organized as follows. Section 2 introduces the background of code translation and motivating examples. Section 3 presents the design of our empirical study, including the research questions, studied translation strategies, translation tasks, LLMs, and evaluation metrics. Section 4 analyzes the experimental results to answer our research questions in detail and discusses the advantages and limitations of pseudocode-based code translation through case studies. Section 6 discusses related work. Finally, Section 7 concludes the paper.
@@ 2.1. Code Translation
Code translation refers to the process of converting programs written in one programming language (PL) into another PL while preserving code semantics (Nguyen et al., 2013; Rozière et al., 2020; Pan et al., 2024). Formally, given an original program $S=\langle s_{1},s_{2},\dots,s_{n}\rangle$ written in the source PL $L_{s}$, code translation aims to produce a translated program $T=\langle t_{1},t_{2},\dots,t_{m}\rangle$ in the target PL $L_{t}$, such that $T$ performs the same functionality as $S$, where $s_{i}$ and $t_{j}$ represent statements in $L_{s}$ and $L_{t}$, respectively. For example, we can translate the C++ program in Figure 1(a) and the Python program in Figure 2(a) to their semantically equivalent Rust program in Figure 1(d) and Java program in Figure 2(d), respectively. In code translation, there may not always be equivalent constructs, statements, and APIs in the target PL for those in the source PL, and vice versa, which makes code translation challenging.
@@
In the past decade, automated code translation approaches have developed from traditional rule-based and statistical machine translation (Nguyen et al., 2013) to neural-model-based methods to automatically learn diverse and complex translation patterns across PLs (Chen et al., 2018; Rozière et al., 2020, 2022; Liu et al., 2023; Jiao et al., 2023). Recently, LLMs pre-trained on massive code corpora demonstrate superior capabilities in code translation (Yang et al., 2024; Ibrahimzada et al., 2025; Pan et al., 2024). These works mainly focus on a code-to-code translation approach, i.e., directly converting the original program in $L_{s}$ to the translated program in $L_{t}$ in an integrated step. However, LLMs may struggle to precisely capture and replicate the semantics of the original program in another PL in a single step, leading to incorrect translations (Pan et al., 2024). To mitigate the gap between source and target PLs, Macedo et al. (2025) explored transitive translation via an intermediate PL, but they still focus on code-to-code translation and fail to handle certain translation tasks. In this study, we explore an alternative code translation approach by emulating human semantic translation in natural language translation (Newmark, 1981), which explicitly conducts code understanding by generating pseudocode as an intermediate step to interpret code intent and logic to facilitate translation.
In code translation, the existence of original programs provides a reference to validate the translated programs. Unlike code generation tasks that rely on manually defined test cases or human evaluation to assess the generated code (Chen et al., 2021; Austin et al., 2021), code translation allows automated correctness assessment by comparing the outputs of the translated and original programs on the same test cases. These enable multiple attempts to be a feasible and meaningful practice to enhance code translation accuracy (Macedo et al., 2025; Zheng et al., 2023). In this study, we explore the translation setup of 10 attempts, which is a typical cost-effective setup that balances cost and gain (Macedo et al., 2025; Wu et al., 2025). It also enables the investigation of combining results from both direct and pseudocode-based code translation (Section 3.3).
@@ 2.2. Motivating Examples
To guide LLMs to solve complicated coding tasks, researchers often decompose a complex task into multiple simpler sub-tasks following human experience (e.g., problem solving/planning and implementation for code generation (Dong et al., 2024; Li et al., 2025; Wu et al., 2025)). Inspired by this, we explore whether decomposing code translation into sub-tasks following human practices can also facilitate LLMs to produce more accurate translations by emulating the successful human workflow.
Recalling the practices in natural language translation, humans typically employ two approaches. When expressions or sentence structures in the source and target languages are sufficiently similar, a literal word-by-word translation is often adequate. However, when literal translation is challenged by significant differences in grammar, idiomatic usage, or context between the source and target languages, humans tend to adopt semantic translation (Newmark, 1981), where they first interpret the underlying meaning of the source sentence and then render it in the target language in the manner best to convey the intended message.
Interestingly, we observe that semantic interpretation also benefits LLMs in code translation. Specifically, we prompt LLMs to explicitly conduct semantic translation via pseudocode-based transitive translation, i.e., first producing a semantic interpretation (pseudocode) for code intent and logic of the original program, and then generating the translated program based on pseudocode. We adopt pseudocode to represent semantics since it is a widely adopted form for describing code intent and logic of algorithms in textbooks and research papers. It also proves effective in guiding LLMs to replicate the code intent and logic in varying PLs and can be synthesized from programs (Wu et al., 2025).
@@
These motivating examples suggest that incorporating pseudocode as a semantic representation for transitive translation can promote accurate code translations, mirroring the benefits of humans using semantic translation for natural languages. This is reasonable as the approach guides LLMs to understand code intent and logic and then perform code generation, decomposing a complex task into simpler sub-tasks, and LLMs are skillful in both tasks (Geng et al., 2024; Jain et al., 2025). This enables a new access to achieving code translation, which may bypass LLMs’ struggle with direct translation.
Despite these benefits, the proper use of pseudocode in code translation remains underexplored, e.g., how to effectively translate code via pseudocode, how effective it is when handling distinct source-target PL pairs, and what strengths it entails. In addition, pseudocode-based code translation may also face challenges, e.g., the potential error propagation from incorrect pseudocode and the inherent ambiguity in natural language. The limitations it entails and potential remedies to overcome the limitations also remain unclear. To bridge these gaps, we conduct a systematic investigation of pseudocode-guided code translation. Our study aims to provide deeper insights into when and how pseudocode can effectively enhance code translation.

### v3

@@ Abstract.
Although large language models (LLMs) show promising potential in code translation, they still struggle to generate accurate translations using the commonly adopted direct code-to-code translation approach, which converts an original program into the target programming language (PL) in a single step. Inspired by the success of incorporating intermediate steps to guide LLMs in resolving challenging tasks, in this study, we explore pseudocode-based code translation. This approach emulates human semantic translation by first interpreting the original program’s intent and logic into pseudocode and then implementing it in the target PL. To understand the effectiveness of this underexplored approach, we present a systematic empirical study on pseudocode-based code translation, aiming to investigate its helpfulness in enhancing the direct translation approach, illuminate its effective usage, and identify its limitations. By comparing direct and pseudocode-based translation on 9,690 translation tasks across six PLs with five popular LLMs, we found that pseudocode-based translation can effectively complement direct translation, particularly when translating from flexible to rigid PLs and handling a low-training-resource PL. Based on the findings, we suggest combining the translation results of both approaches for test-based selection to leverage their complementary strengths. We also reveal the advantages of pseudocode-based translation in decoupling the code understanding and generation burden on complicated programs and mitigating distractions from PL-specific implementations in original programs, as well as its limitations due to incorrect, incomplete, or ambiguous pseudocode. Our study sheds light on the effective use of pseudocode-based translation and provides evidence to help enhance LLMs in code translation.
@@
copyright: ccjournal: TOSEMjournalyear: 2026journalvolume: 1journalnumber: 1publicationmonth: 1doi: 10.1145/3790101ccs: General and reference Empirical studiesccs: Software and its engineering Software maintenance tools
@@ 1. Introduction
Code translation, also known as transpilation, refers to the automatic conversion of a program written in one programming language (PL) to another while preserving its behaviors (Nguyen et al., 2013; Rozière et al., 2020; Pan et al., 2024). With the rapid evolution of PLs and the diverse requirements of software applications, code translation has gained significant attention in both research and industry due to its wide range of practical applications. For example, code translation facilitates the migration of legacy software systems built based on obsolete PLs to modern PLs (Echeverria et al., 2015; Kalia et al., 2021a, b). It also facilitates efficient prototyping and development across multiple PLs for multi-platform software (Liu et al., 2023; Nguyen et al., 2015; Wu et al., 2010). However, achieving accurate automated code translation remains a challenging task due to the inherent differences in syntax, features, and APIs across various PLs (Pan et al., 2024).
@@
Over the past few decades, various approaches have been proposed to automate code translation and address the associated challenges (Chen et al., 2025a; Jiao et al., 2023). Earlier methods primarily relied on statistical or neural machine translation (Nguyen et al., 2013; Chen et al., 2018; Rozière et al., 2020, 2022; Liu et al., 2023), while recent practices start leveraging the powerful large language models (LLMs) as code translators (Pan et al., 2024; Yang et al., 2024). However, the state-of-the-art practices with LLMs mainly use a direct code-to-code translation approach, where the original program is taken as input to generate the translated program in a single step (Yang et al., 2024). In this manner, LLMs may perform multiple tasks implicitly, such as understanding the semantics of the original program and implementing the semantics into the target program in another PL. Studies have shown that such direct translation remains challenging, notably across PLs with significant differences (Pan et al., 2024; Macedo et al., 2025).
@@
Decomposing challenging tasks into intermediate steps has proven effective in guiding LLMs to emulate successful human workflows in various coding tasks (e.g., splitting planning and implementation for code generation (Dong et al., 2024; Li et al., 2025) and fault localization and patch generation for debugging (Xia et al., 2024)). However, there are few experiences about whether decomposition also helps generate accurate code translations. Notably, we observe that code translation can also benefit from explicitly implementing intermediate steps that emulate human practices in translation. Specifically, human translators often perform semantic translation (Newmark, 1981) based on the semantics of the given text, first interpreting its meaning and then rendering the meaning in the target language (Newmark, 1981). In addition, a recent study demonstrates that pseudocode, as a PL-agnostic representation of code intent and logic widely used in textbooks and research papers, can effectively guide code generation across PLs (Wu et al., 2025). Based on these insights, we realize that emulating semantic translation for code based on pseudocode can be a solution for the tasks that direct translation struggles with. For example, when translating a simple C++ program (Figure 1(a)) to Rust, all five studied LLMs failed to produce an accurate translation via a one-step direct translation (Figure 1(b)). Meanwhile, following a clear pseudocode (Figure 1(c)) summarized from the original C++ program, these LLMs can successfully implement the original program’s functionality in Rust (Figure 1(d)). Nevertheless, the effectiveness, advantages, and limitations of the promising pseudocode-based code translation approach compared to the commonly adopted direct translation remain unclear.
@@
In this work, we conduct the first systematic empirical study on pseudocode-based code translation to bridge the gap in understanding this under-explored approach. Specifically, we assessed the performance of five translation strategies: the widely-adopted direct translation approach, two pseudocode-based strategies, and two hybrid strategies that combine direct and pseudocode-based translation. Through four research questions (RQs), we investigate the overall effectiveness of pseudocode-based code translation compared with the widely-adopted direct translation (RQ1), its helpfulness for varying PL pairs (RQ2), the advantages of pseudocode as an intermediary compared to a concrete PL, which also proved helpful in code translation (Macedo et al., 2025) (RQ3), and the impact of pseudocode quality on translation performance (RQ4). We also examine the successes and failures of pseudocode-based code translation via case studies and discuss future directions for further harnessing its potential benefits. The study is conducted with solution programs in six popular PLs for 323 LeetCode problems in three difficulty levels in LiveCodeBench (Jain et al., 2025), resulting in 9,690 translation tasks (i.e., 6 source PLs $\times$ 5 target PLs $\times$ 323 problems), for five popular LLMs.
@@
Our experimental results reveal several interesting findings and actionable insights. Specifically, we identify the complementary role of pseudocode-based code translation in effectively handling translation tasks that direct translation struggles with. By combining the results of direct translation and pseudocode-based translation, the studied LLMs achieve average improvements in translation accuracy over direct translation alone by 4.10%, 7.44%, and 13.76% on easy-, medium-, and hard-level problems, respectively. We also demonstrate that pseudocode-based translation can benefit various source-target PL pairs, with more significant benefits for translation from flexible PLs (e.g., Python) to more rigid PLs (e.g., Go), or involving Rust PL that has limited training resources. These findings suggest the adoption of a hybrid strategy that combines direct and pseudocode-based translation results with further test-based selection to leverage the complementary strengths of both approaches, particularly when handling the highly benefited PLs/PL pairs. Moreover, with higher-quality pseudocode, we observed that the studied LLMs can achieve pass rates of 0.9646–0.9835, 0.8861–0.9512, and 0.6747–0.8286 on three-level tasks, respectively, identifying bottlenecks of pseudocode-based code translation in both code understanding and code generation capabilities of LLMs. Our case studies further illustrate the advantages of pseudocode-based translation in decoupling the code understanding and generation burdens on complicated programs and mitigating distractions from the PL-specific details in the original programs, as well as its limitations caused by incorrect, incomplete, or ambiguous pseudocode. These findings illuminate the effective adoption of pseudocode-based translation to enhance LLMs’ code translation accuracy and provide evidence for future research on improving LLM-driven code translation.
@@
Figure 1. An Example of Correct C++-to-Rust Translation by Qwen2.5-Coder-32B-Instruct based on Pseudocode, where Direct Translation Failed. (The other four studied LLMs show similar symptoms in this case.)
@@
  • We conduct the first systematic empirical study on pseudocode-based code translation, exploring the effectiveness, advantages, and limitations of emulating human semantic code translation practices via pseudocode in improving code translation accuracy of LLMs.
  • We compare four pseudocode-based translation strategies with direct translation across 9,690 translation tasks involving six popular PLs. With five popular LLMs as the translator, we systematically investigate the effectiveness of pseudocode-based code translation in enhancing direct translation, its helpfulness for varying PL pairs, and its advantages and limitations, aiming to provide insights for its effective adoption and evidence for its future enhancement.
  • Our study reveals intriguing findings, including the complementary role of pseudocode-based translation in enhancing direct code translation, the highly benefited PLs/PL pairs, and the potential and bottlenecks of pseudocode-based code translation. These findings yield actionable insights, including the adoption of a hybrid strategy that combines direct and pseudocode-based translation results to enhance code translation accuracy, as well as identify improvement space in the generation of high-quality pseudocode to further harness the benefits of the approach.
We release our experimental data, including 1,938 solution programs in six popular PLs collected for 323 LeetCode problems in LiveCodeBench-v5 to translate, as well as scripts to conduct direct and pseudocode-based translation and evaluate multilingual translation results. We will open-source them in our artifact (Artifact of this paper, ) to facilitate future research.
The rest of the paper is organized as follows. Section 2 introduces the background of code translation and motivating examples. Section 3 presents the design of our empirical study, including the research questions, studied code translation strategies, tasks, LLMs, and evaluation metrics. Section 4 analyzes the experimental results to answer our research questions in detail and discusses the advantages and limitations of pseudocode-based code translation through case studies. Section 5 and Section 6 discuss the threats to validity and related work, respectively. Finally, Section 7 concludes the paper.
@@ 2.1. Code Translation
Code translation refers to the process of converting programs written in a programming language (PL) into another while preserving their code semantics (Nguyen et al., 2013; Rozière et al., 2020; Pan et al., 2024). Formally, given an original program $S=\langle s_{1},s_{2},\dots,s_{n}\rangle$ written in the source PL $L_{s}$, code translation aims to produce a translated program $T=\langle t_{1},t_{2},\dots,t_{m}\rangle$ in the target PL $L_{t}$, such that $T$ performs the same functionality as $S$, where $s_{i}$ and $t_{j}$ represent statements in $L_{s}$ and $L_{t}$, respectively. For example, we can translate the C++ program in Figure 1(a) and the Python program in Figure 2(a) to their semantically equivalent Rust program in Figure 1(d) and Java program in Figure 2(d), respectively. In code translation, there may not always be equivalent constructs, statements, and APIs in the target PL for those in the source PL, and vice versa, which makes code translation challenging (Yang et al., 2024; Macedo et al., 2025).
@@
In the past decade, automated code translation approaches have developed from traditional rule-based and statistical machine translation (Nguyen et al., 2013) to neural-model-based methods to automatically learn diverse and complex translation patterns across PLs (Chen et al., 2018; Rozière et al., 2020, 2022; Liu et al., 2023; Jiao et al., 2023). Recently, LLMs pre-trained on massive code corpora demonstrate superior capabilities in code translation (Yang et al., 2024; Ibrahimzada et al., 2025; Pan et al., 2024). These works mainly focus on a direct one-step code-to-code translation approach, i.e., directly converting the original program in $L_{s}$ to the translated program in $L_{t}$ in an integrated step. However, LLMs may struggle to precisely capture and replicate the semantics of the original program in another PL in a single step, leading to incorrect translations (Pan et al., 2024). To mitigate the gap between source and target PLs, Macedo et al. (2025) explored transitive translation via an intermediate PL, i.e., chaining two code-to-code translations. But they still fail to handle certain translation tasks. In this study, we explore an alternative code translation approach by emulating human semantic translation in natural language translation (Newmark, 1981), which explicitly conducts code interpretation by generating pseudocode as an intermediate step to understand code intent and logic to facilitate translation.
A property of the code translation task is that the original program is available, enabling automatic correctness assessment of multiple translation candidates. Specifically, unlike the code generation tasks where the result correctness is typically evaluated using manually prepared test assertions or human judgment (Chen et al., 2021; Austin et al., 2021), code translation allows automatic correctness assessment of the translation results by comparing the outputs of the translated and original programs on the same test inputs. This property makes it practical and meaningful to generate multiple translation candidates to increase the likelihood of getting a correctly translated program, since the original program can help identify the best candidate as the final result via automatic correctness assessment (Rozière et al., 2020; Macedo et al., 2025; Zheng et al., 2023). In this study, we adopt ten generation attempts, which represents a common cost-effective trade-off between computational cost and performance gains on coding tasks (Macedo et al., 2025; Wu et al., 2025). The multiple-attempt setup also allows us to easily explore the effectiveness of combining results from both direct and pseudocode-based translation approaches (Section 3.3).
@@ 2.2. Motivating Examples
To guide LLMs to solve complicated coding tasks, researchers propose to decompose a complex task into multiple simpler sub-tasks following human experience (e.g., problem solving/planning and implementation for code generation (Dong et al., 2024; Li et al., 2025; Wu et al., 2025)). Inspired by this, we explore whether decomposing code translation into sub-tasks following human practices can also facilitate LLMs to produce more accurate code translations by emulating the successful human workflow.
Recall the human practices in natural language translation. When the expressions or sentence structures in source and target languages are sufficiently similar, a literal word-by-word translation is often feasible. Meanwhile, when human translators struggle with the significant differences in grammar, idiomatic usage, or context between source and target languages, humans tend to adopt semantic translation (Newmark, 1981), where they first interpret the underlying meaning of the source sentence and then render it in the target language in the manner best to convey the intended message.
Interestingly, we observe that semantic interpretation also benefits LLMs in code translation. Specifically, we prompt LLMs to explicitly conduct semantic translation via pseudocode-based transitive translation, i.e., first producing a semantic interpretation (pseudocode) for code intent and logic of the original program, and then generating the translated program based on pseudocode. We use pseudocode to represent semantics because it is a common form for describing code intent and logic of algorithms in textbooks and research papers. Also, it proves effective in guiding LLMs to replicate the code intent and logic in varying PLs and can be synthesized from programs (Wu et al., 2025).
@@
These motivating examples suggest that incorporating pseudocode as a semantic representation and conducting transitive translation can promote accurate code translations, mirroring the benefits of humans using semantic translation for natural languages. This is reasonable as the approach guides LLMs to understand code intent and logic and then perform code generation, decomposing a complex task into simpler sub-tasks, and LLMs are skillful in both tasks (Geng et al., 2024; Jain et al., 2025). This enables a new access to achieving code translation, which may bypass LLMs’ struggle with direct translation.
Despite these benefits, the proper use of pseudocode in code translation remains underexplored, e.g., how to effectively translate code via pseudocode, how effective it is when handling different source-target PL pairs, and what strengths it entails. In addition, pseudocode-based code translation may also face challenges, e.g., the potential error propagation from incorrect pseudocode and the inherent ambiguity in natural language (Liu et al., ). The limitations it entails and potential remedies to overcome the limitations also remain unknown. To bridge these gaps, we conduct a systematic investigation of pseudocode-guided code translation. Our study aims to provide insights into how emulating semantic translation via pseudocode can effectively enhance code translation.